# PaTAS: A Parallel System for Trust Propagation in Neural Networks Using Subjective Logic

### v2

@@ 0 Trusted Trusted Trusted
PaTAS provides a principled mechanism for evaluating the runtime trustworthiness of neural network outputs. Beyond dynamic inference assessment, it can also estimate a static trust opinion of the network itself. The core principle is that a trustworthy model should preserve trust: a fully trusted input producing a highly trusted output indicates that the model does not erode trust during inference. Accordingly, the overall model trustworthiness can be defined as the trust score of the PaTAS feedforward function ($PaTAS_{FF}$) under a fully trusted input:
@@
which quantifies how well trust is maintained throughout the network. Using this definition, we compared the Inference Path Trust Assessment (IPTA) for benign and adversarial inputs in Experiment VI-A3, where the adversarial case used a $4\times 4$ patch injection. Results show a clear degradation in trust for adversarial samples: benign inputs achieved $t=0.484,d=0,u=0.516$, while patched inputs dropped to $t=0.449,d=0,u=0.551$. This demonstrates the sensitivity of PaTAS to local perturbations, offering a quantitative indicator of reduced reliability.
Although trust mass and accuracy are often correlated, our results reveal meaningful divergences. In the breast cancer task (Table II), assigning a mixed trust profile $(0.25,0,0.75)$ to all features and labels yields a final trust mass of $0.38$, higher than the fully uncertain-features case, yet the latter achieves better test accuracy (96.49% vs. 90%). Thus, higher trust mass does not always imply superior predictive performance. A similar effect appears in the poisoned MNIST experiments (Table IV): the clean digit “6” attains slightly higher accuracy than “3” (93.52% vs. 93.14%), but IPTA assigns higher trust to “3” ($t=0.484$) than to “6” ($t=0.482$). These differences show that PaTAS captures aspects of reliability beyond accuracy, reflecting the stability and interpretability of inference paths.
@@
In practice, precise dataset trustworthiness estimates may be unavailable. PaTAS can still operate by initializing dataset trust to a fully uncertain (vacuous) opinion. While this limits the use of prior trust evidence, it preserves the capacity to propagate input-level trust through the model during inference, supporting meaningful trust evaluation even without explicit dataset provenance.
@@ VII-B Trust Assessment and AI Security Threats
Figure 7: Attack surface in the AI lifecycle pipeline, from data collection to model deployment.
AI systems are exposed to diverse attacks targeting different stages of the lifecycle. Fig. 7 outlines six key phases: data collection, cleaning and labeling, dataset assembly, network design, model training, and deployment for inference.
@@ VIII Conclusion
This paper presented the Parallel Trust Assessment System (PaTAS), a framework that models and propagates trust in neural networks through Subjective Logic. PaTAS introduces a parallel computational structure based on Trust Nodes and Trust Functions, enabling principled trust propagation alongside standard feedforward and backpropagation processes. The proposed Parameter Trust Update and Inference-Path Trust Assessment (IPTA) mechanisms jointly quantify how input quality, learned parameters, and activation paths contribute to the overall trustworthiness of model predictions.
@@
Appendix A Subjective Logic Operators and Symbols
@@
Appendix B PaTAS Operations
(a) Trust Aggregation Process
(b) Example of Trust Aggregation Process
Figure 8: Output-Trust Aggregation
Fig. 8(a) illustrates this process. Each Trust Node in the output layer generates a trust score $T_{y_{1}},T_{y_{2}},\ldots,T_{y_{k}}$, corresponding to the network’s output components. These values are then passed to an aggregation mechanism that produces a unified output trust value $T_{y}$, formally defined as:
$T_{y}=\text{Agg}\left((T_{y_{j}})_{j\in[1,k]}\right)$
where Agg is an aggregation function based on the reasoning operations defined in the PaTAS.
Fig. 8(b) presents a conceptual example in which a virtual observer $A$ receives individual trust from each output neuron. Observer $A$ integrates these trust scores using the aggregation function to arrive at a single opinion on the network’s overall prediction trustworthiness.
The aggregation mechanism can be adapted to different application requirements, such as majority fusion, weighted fusion, or uncertainty-sensitive operators, depending on how the output trust should reflect the underlying trust semantics.
Appendix C Theorems Proofs
C-A Proof for Theorem 1
@@
C-B Proof for Theorem 2
@@ Appendix D Detailed results
All the results are depicted in Figs. 11, 12, 9, 16, 18, 19, 13, 17, 14 and 15.
(a) 5 Hidden neurons
(b) 10 Hidden neurons
(c) 20 Hidden neurons
(d) 20 Hidden neurons with Fully Trusted Assessment
Figure 9: MNIST with Vacuous Trust Assessment
(a) 5 Hidden neurons
(b) 10 Hidden neurons
(c) 20 Hidden neurons
(d) 20 Hidden neurons with Fully Trusted Assessment
Figure 10: MNIST with Vacuous Trust Assessment
(a) xdistrust+ydistrust
(b) xdistrust+yvacuous
(c) xdistrust+ytrust
(d) xvacuous+ydistrust
(e) xvacuous+yvacuous
(f) xvacuous+ytrust
(g) xtrust+ydistrust
(h) xtrust+yvacuous
(i) xtrust+ytrust
Figure 11: Cancer Model with epsilon = 0.01
(a) xdistrust+ydistrust
(b) xdistrust+yvacuous
(c) xdistrust+ytrust
(d) xvacuous+ydistrust
(e) xvacuous+yvacuous
(f) xvacuous+ytrust
(g) xtrust+ydistrust
(h) xtrust+yvacuous
(i) xtrust+ytrust
Figure 12: Cancer Model with epsilon = 0.1
(a) (0.25, 0, 0.75)
(b) (0.25, 0.25, 0.5)
Figure 13: Degradation
@@
(a) 1 pixel
(b) 4×4 pixels
(c) 20×20 pixels
(d) 27×27 pixels
Figure 16: MNIST poisoned 20 hidden neurons
@@
(a) 1 pixel
(b) 4×4 pixels
(c) 20×20 pixels
(d) 27×27 pixels
Figure 18: MNIST poisoned 10 hidden neurons

### v3

@@ 0 Trusted Trusted Trusted
PaTAS is a framework for evaluating the runtime trustworthiness of neural network outputs. Beyond dynamic inference assessment, it can also estimate a static trust opinion of the network itself. The core principle is that a trustworthy model should preserve trust: a fully trusted input producing a highly trusted output indicates that the model does not erode trust during inference. Accordingly, the overall model trustworthiness can be defined as the trust score of the PaTAS feedforward function ($PaTAS_{FF}$) under a fully trusted input:
@@
which quantifies how well trust is maintained throughout the network. Using this definition, we compared the Inference Path Trust Assessment (IPTA) for benign and adversarial inputs in Experiment VI-A3, where the adversarial case used a $4\times 4$ patch injection. Results show a degradation in trust for adversarial samples: benign inputs achieved $t=0.484,d=0,u=0.516$, while patched inputs dropped to $t=0.449,d=0,u=0.551$. These observations demonstrate the sensitivity of PaTAS to local perturbations, offering a quantitative indicator of reduced reliability.
Although trust mass and accuracy are often correlated, our results reveal meaningful divergences. In the breast cancer task (Table II), assigning a mixed trust profile $(0.25,0,0.75)$ to all features and labels yields a final trust mass of $0.38$, higher than the fully uncertain-features case, yet the latter achieves better test accuracy (96.49% vs. 90%). Thus, higher trust mass does not always imply superior predictive performance. A similar effect appears in the poisoned MNIST experiments (Table IV): the clean digit “6” attains slightly higher accuracy than “3” (93.52% vs. 93.14%), but IPTA assigns higher trust to “3” ($t=0.484$) than to “6” ($t=0.482$) which was trained with partially poisoned training data.
These differences reflect that PaTAS evaluates the reliability of the inference path, not just predictive accuracy. The digit “6” was a poisoned class during training, so the parameters involved in predicting “6” are less trusted than those for “3,” even though the test samples are clean.
@@
In practice, precise dataset trustworthiness estimates may be unavailable, especially for large training datasets where the provenance of individual samples is difficult to establish. PaTAS can still operate in such cases by initializing dataset trust to a fully uncertain (vacuous) opinion. While this limits the use of prior trust evidence during training, it preserves the ability to propagate trust assessments at inference time. For a specific query, the circumstances under which the input features were collected are often easier to assess, making input-level trust more practical to obtain than dataset-level trust. PaTAS can therefore support meaningful trust evaluation even when explicit dataset provenance is missing. Likewise, if trust in a particular input cannot be assessed, a fully uncertain opinion may be used as a fallback.
@@ VII-B Trust Assessment and AI Security Threats
AI systems are exposed to diverse attacks targeting different stages of the lifecycle. We can distinguish six key phases: data collection, cleaning and labeling, dataset assembly, network design, model training, and deployment for inference.
@@ VIII Conclusion
This paper presented the Parallel Trust Assessment System (PaTAS), a framework that models and propagates trust in neural networks through Subjective Logic. PaTAS introduces a parallel computational structure based on Trust Nodes and Trust Functions, enabling trust propagation alongside standard feedforward and backpropagation processes. The proposed Parameter Trust Update and Inference-Path Trust Assessment (IPTA) mechanisms jointly quantify how input quality, learned parameters, and activation paths contribute to the overall trustworthiness of model predictions.
@@
Several directions remain open for further investigation. First, the effect of activation functions on trust propagation has not been studied in detail. In our experiments we used ReLU, which has convenient properties for trust reasoning. Other activation functions, for example sigmoid, tanh, or GELU, may influence how trust is transformed across layers. It is also necessary to explore the impact of various in-between layers in neural networks, such as convolutional or transformer layers, on the PaTAS’ trust evaluation. Understanding these effects is an important step toward applying PaTAS in a broader range of architectures.
Second, the computational cost of PaTAS deserves a more systematic evaluation. PaTAS operates in parallel with the neural network, and this parallelism can be exploited to optimize both inference and the construction of trust assessments. Benchmarking runtime and memory consumption on larger models will help identify opportunities for further optimization.
Third, a quantitative comparison with DeepTrust and related trust assessment frameworks is planned. This comparison will help clarify conceptual differences and highlight cases where PaTAS provides advantages in terms of interpretability or performance.
Appendix A Subjective Logic Operators and Notation for Parameter-Trust Update
@@
TABLE VII: Symbols Used in the Parameter-Trust Update Subsection and Algorithm
Symbol Meaning / Description
Inputs to the Algorithm
$g$ Collection of gradients for all parameters in the batch
$T_{y}$ Trust opinion on label $y$
$\epsilon$ Gradient sensitivity threshold in NodeTrust
Batch and Layer Quantities
$T_{y_{\text{batch}}}$ Aggregated trust over labels in the current batch
$l$ Layer index
$n_{i}^{(l)}$ Neuron $i$ in layer $l$
$\mathcal{N}(i)$ Set of incoming edges to neuron $i$
Gradient Evidence
$g^{(l)}_{i,j}$ Gradient of loss w.r.t. $\theta^{(l)}_{i,j}$
$g^{(l)}_{i}$ Gradient vector for neuron $n_{i}^{(l)}$
$r$ Count of weak gradients: $|g^{(l)}_{i,j}|<\epsilon$
$s$ Count of strong gradients: $|g^{(l)}_{i,j}|\geq\epsilon$
Trust Values Computed in the Algorithm
$T_{n_{i}\mid y_{\text{batch}}}$ Trust in neuron $n_{i}$ conditioned on batch labels
$T_{n_{i}\mid\overline{y_{\text{batch}}}}$ Trust in neuron under incorrect labels (vacuous)
$T_{n_{i}\parallel Y_{\text{batch}}}$ Deduced trust in neuron $n_{i}$
Parameter Trust
$T_{\theta^{(l)}_{i,j}}$ Trust opinion on parameter $\theta^{(l)}_{i,j}$
$T_{lr}$ Trust in the learning rate
$T_{x^{(l-1)}_{j}}$ Trust in the input feature to parameter $\theta^{(l)}_{i,j}$
Operators Used
$\bigwedge$ Batch-wise fusion of trust opinions
[…]
$\odot$ SL binomial multiplication
[…]
Appendix B Symbols and Parameters Used in PaTAS
TABLE VIII: List of Symbols and Parameters Used in this work
Symbol Meaning / Description
Neural Network Quantities
$x$ Input feature vector
$y$ Ground-truth label
$y^{\prime}=f_{\Theta}(x)$ Output of the neural network
$\Theta=(W,b,f)$ Neural-network parameters (weights, biases, activations)
$W^{(l)}$ Weight matrix at layer $l$
$b^{(l)}$ Bias vector at layer $l$
$f^{(l)}(\cdot)$ Activation function at layer $l$
$z^{(l)}$ Pre-activation vector of layer $l$
$x^{(l)}$ Activation vector of layer $l$
$n_{i}^{(l)}$ Neuron $i$ in layer $l$
$\theta^{(l)}_{i,j}$ Parameter from neuron $j$ in layer l1l\!-\!1 to neuron $i$ in layer $l$
$\delta^{(l)}_{i}$ Backpropagated error signal of neuron $i$ in layer $l$
$x^{(l-1)}_{j}$ Activation of neuron $j$ in the previous layer
$\mathcal{N}(i)$ Set of incoming neighbors of neuron $i$
Gradients and Learning Dynamics
$g$ Collection of gradients for all network parameters
$g^{(l)}_{i,j}$ Gradient w.r.t. weight $\theta^{(l)}_{i,j}$
$g^{(l)}_{i}$ Vector of gradients for all incoming parameters of neuron $n_{i}^{(l)}$
$l_{r}$ (or $\mathrm{lr}$) Learning rate
$\mathcal{L}$ Loss function
Trust Assessments and Trust Nodes
$T(x)$ Trust assessment of input features
$T_{y}$ Trust opinion on label $y$
$T_{y_{\text{batch}}}$ Aggregated trust opinion over all labels in a batch
$T_{\theta^{(l)}_{i,j}}$ Trust opinion on parameter $\theta^{(l)}_{i,j}$
$T_{x^{(l)}_{i}}$ Trust opinion on activation $x^{(l)}_{i}$
$T_{x^{(l)}_{j}}$ Trust opinion of feature contributing to edge $(j\to i)$
$T_{lr}$ Trust opinion on the learning rate
$T_{n_{i}\mid y_{\text{batch}}}$ Trust in neuron $n_{i}$ conditioned on batch labels
$T_{n_{i}\mid\overline{y_{\text{batch}}}}$ Trust in neuron given incorrect labels (vacuous opinion)
$T_{n_{i}\parallel Y_{\text{batch}}}$ Deduced trust in neuron $n_{i}$ after combining conditional evidence
Subjective Logic Opinions and Evidence
$\omega=(b,d,u,a)$ SL binomial opinion: belief, disbelief, uncertainty, base rate
$r$ Positive evidence count (from weak gradients)
$s$ Negative evidence count (from strong gradients)
$\epsilon$ Gradient sensitivity threshold in NodeTrust
Trust Operators (SL Operators + PaTAS-specific)
$\oplus$ SL fusion operator (cumulative or averaging fusion)
$\otimes$ SL trust-discounting operator
$\ominus$ SL opinion-revision operator
$\odot$ SL binomial multiplication
[…]
$\bigwedge$ Batch-wise fusion of trust opinions
PaTAS System Components
$\mathrm{Pf}_{\Theta}$ Parallel Trust Function (PaTAS trust feedforward)
TNN Trust Nodes Network (parallel trust architecture)
GenIPTA Generator of the Inference-Path Trust Assessment
IPTA Inference-Path Trust Assessment for a single inference
Appendix C Theorems and Proofs (Section V-D)
C-A Key theorems
Theorem 4 (Convergence of Subjective Logic Arithmetic Sequence).
Let $(\Omega,\ominus)$ be a group. $\Omega$ is a set of opinions specified as $0,1]^{4}$ Let $\omega_{n}=(b_{n},d_{n},u_{n},a_{n})\in\Omega$ be a sequence defined recursively by the operator $\ominus$ as follows:
$\omega_{n+1}=\omega_{n}\ominus q,$
where $q\in\Omega$ and the operator $\ominus$ is a fusion operator.
Then the sequence $(\omega_{n})$ converges in $0,1]^{4}$ to a limit:
$\begin{cases}\omega_{0},&\text{if }q=(0,0,1,a_{0}),\\ q&\text{otherwise}.\end{cases}$
@@ C-A Key theorems
Theorem 5 (Convergence of Subjective Logic Geometric Sequence).
Let $(\Omega,\odot)$ be a group defined as in Theorem 4. Let $\omega_{n}=(b_{n},d_{n},u_{n},a_{n})\in\Omega$ be a sequence defined recursively by the operator $\odot$ as follows:
$\omega_{n+1}=\omega_{n}\odot q,$
  • $a_{n}=a_{0}a_{q}^{n}\to 0$ as $n\to\infty$,
  • $d_{n}\to 1$ if $d_{q}>0$, and $d_{n}=d_{0}$ if $d_{q}=0$,
  • $b_{n}\to 0$ if $p_{q}=b_{q}+a_{q}u_{q}<1$, and $b_{n}=b_{0}$ if $p_{q}=1$.
@@ 2. Convergence of $d_{n}$: The recurrence relation is:
C-B Proof for Theorem 1
Proof.
Moreover, the stability of $T_{x}$, $T_{y}$, and $T_{l_{r}}$ implies that trust inputs to the update mechanism are stable. Consequently, each new trust update for $T_{\theta}$ is computed using consistent and bounded evidence, which over time results in the stabilization of the subjective opinions assigned to $T_{\theta}$.
In detail:
  • $g$ converges so $T_{\theta\mid y_{\text{batch}}}$ will remain the same.
  • since $T_{\theta\mid\mid{Y_{\text{batch}}}}$ is a deterministic calculation from the convergent terms $T_{\theta\mid y_{\text{batch}}}$ and $T_{y_{batch}}$, it follows that $T_{\theta\mid\mid{Y_{\text{batch}}}}$ also converges
[…]
  • The function $f_{upd}$ in our implementation is based on $T_{\theta}$, $T_{l_{r}}$ and $T_{x}$} which all converge.
@@ Appendix D Detailed results
All the results are depicted in Figures in Sections D-A, D-B, D-C, D-D, D-E, D-F, D-G and D-H.
D-A Plots for Cancer Model with $\epsilon=0.01$ (Section VI-A1)
Figure 8: xdistrust+ydistrust
Figure 9: xdistrust+yvacuous
Figure 10: xdistrust+ytrust
Figure 11: xvacuous+ydistrust
Figure 12: xvacuous+yvacuous
Figure 13: xvacuous+ytrust
Figure 14: xtrust+ydistrust
Figure 15: xtrust+yvacuous
Figure 16: xtrust+ytrust
D-B Cancer Model with $\epsilon=0.1$(Section VI-A1)
Figure 17: xdistrust+ydistrust
Figure 18: xdistrust+yvacuous
Figure 19: xdistrust+ytrust
Figure 20: xvacuous+ydistrust
Figure 21: xvacuous+yvacuous
Figure 22: xvacuous+ytrust
Figure 23: xtrust+ydistrust
Figure 24: xtrust+yvacuous
Figure 25: xtrust+ytrust
D-C Results for trust Degradation (Section VI-A1)
Figure 26: (0.25, 0, 0.75)
Figure 27: (0.25, 0.25, 0.5)
D-D Accuracy Evolution for the Cancer Model (Section VI-A1)
@@
D-E MNIST with Vacuous Trust Assessment (Section VI-A2)
Figure 30: 5 Hidden neurons
Figure 31: 10 Hidden neurons
Figure 32: 20 Hidden neurons
Figure 33: 20 Hidden neurons with Fully Trusted Assessment
(a) 5 Hidden neurons
(b) 10 Hidden neurons
(c) 20 Hidden neurons
(d) 20 Hidden neurons with Fully Trusted Assessment
Figure 34: MNIST with Vacuous Trust Assessment
D-F MNIST poisoned 20 hidden neurons (Section VI-A3)
Figure 35: 1 pixel
Figure 36: 4×4 pixels
Figure 37: 20×20 pixels
Figure 38: 27×27 pixels
@@ D-F MNIST poisoned 20 hidden neurons (Section VI-A3)
D-G MNIST poisoned 10 hidden neurons
Figure 40: 1 pixel
Figure 41: 4×4 pixels
Figure 42: 20×20 pixels
Figure 43: 27×27 pixels
D-H Example for random dataset trust assessment